OpenStudy (anonymous):

Integral of ydx+x2dy where C=C1+C2. C1 is the path of the straight line segment from the origin, (0,0) to the point (4,28). C2 is the path of the parabola y=−x^2+10x+4 from the point (4,28) to the point (7,25) .

OpenStudy (anonymous):

For C1, I have x=4t, dx=4 dt and y=28t, dy=28 dt So the integral is from 0 to 4: (4t*28)+(28*16t^2) dt For C2, I have x=t, y=-t^2+10t+4 giving dx=dt and dy=-2t+10 So the integral is from 4 to 7: -2t^3+9t^2+10t+4 C1=10453.3 and C2=-58.5 giving C=C1+C2= 10453.3-58.5=10394.8 This is wrong though. Can someone explain what I did wrong?

OpenStudy (anonymous):

@Luigi0210

OpenStudy (luigi0210):

Too advanced for me xD @whpalmer4 @ganeshie8 @mathmale

OpenStudy (anonymous):

Thanks for the help tags luigi

ganeshie8 (ganeshie8):